Suppose that today you deposit the amount of $550 in a bank. How much would the deposit grow to at the end of year 16 if the interest rate is 7.00%?
$

Place your answer in dollars and cents. Do not include a dollar sign or comma in your answer. Work your analysis with at least four decimal places of accuracy.
Answer #1

We know that FV = PV*(1+r)^t

Here, PV = $550

r = 7%

t= 16 year

substituting in formula, we get

FV = 550*(1+0.07)^16 = $1623.6901

the deposit grow to $1623.6901 at the end of year 16 if the interest rate is 7.00%